### v1

@@ Abstract
The Laplace approximation (LA) of the Bayesian posterior is a Gaussian distribution centered at the maximum a posteriori estimate. Its appeal in Bayesian deep learning stems from the ability to quantify uncertainty post-hoc (i.e., after standard network parameter optimization), the ease of sampling from the approximate posterior, and the analytic form of model evidence. However, an important computational bottleneck of LA is the necessary step of calculating and inverting the Hessian matrix of the log posterior. The Hessian may be approximated in a variety of ways, with quality varying with a number of factors including the network, dataset, and inference task. In this paper, we propose an alternative framework that sidesteps Hessian calculation and inversion. The Hessian-free Laplace (HFL) approximation uses curvature of both the log posterior and network prediction to estimate its variance. Only two point estimates are needed: the standard maximum a posteriori parameter and the optimal parameter under a loss regularized by the network prediction. We show that, under standard assumptions of LA in Bayesian deep learning, HFL targets the same variance as LA, and can be efficiently amortized in a pre-trained network. Experiments demonstrate comparable performance to that of exact and approximate Hessians, with excellent coverage for in-between uncertainty.
@@ 1 Introduction
Bayesian neural networks, where a prior is placed on model weights, offer an opportunity to understand uncertainty and direct experimentation. However, exact computation of the model-weight posterior and predictive distributions is prohibitive given size and complexity. It is instead appealing to approximate it using the maximum a posteriori (MAP) solution and the curvature thereat. However, even this curvature, characterized by a Hessian, can be challenging to compute.
@@
$\hat{\theta}$ $=\arg_{\theta}\max\mathcal{L}_{\theta}$ (1)
@@
for some observation likelihood $p(y\;|\;\phi)$ and prior $p(\theta)$. This is equivalent to optimizing a loss (e.g., squared loss for Gaussian likelihood or cross-entropy for categorical likelihood) regularized by $\log p(\theta)$ and means that computing the MAP of a Bayesian neural network is akin to fitting a vanilla neural network.
As detailed in the next section, the Laplace approximation of of the posterior distribution leverages the (negative) Hessian of the log joint distribution at the MAP (Mackay, 1992),
$P=-\nabla\nabla_{\theta}\mathcal{L}_{\hat{\theta}}.$ (3)
Calculating and inverting $P$ are the central steps in performing Laplace approximation. For high dimensional $\theta$ – common in deep neural networks – these steps are a prohibitive computational bottleneck even for an approximate posterior. One remedy is further approximation of the Hessian as well, but these may ignore important directions of curvature.
In this paper we propose Hessian-free Laplace (HFL), which directly approximates the uncertainty of the predicted mean in the Laplace approximation without explicitly computing and inverting the Hessian (while still assuming it exists, as in Laplace). Instead, one additional point estimate is required, the prediction-regularized MAP, derived from the network objective with a small amount of regularization added. Under the assumptions that the Laplace approximation is used, we find that the rescaled difference in network outputs given by the MAP and the prediction-regularized MAP recovers the Laplace variance.
Our contributions in this paper are as follows:
  • We develop a set of novel methods around regularization-based uncertainty that targets the predictive variance of linearized Laplace without explicit Hessian calculations. Instead, variance is derived from the change in the MAP when adjusting the regularizer.
  • In particular, we present Hessian-free Laplace (HFL), a method that estimates the predictive variance of a network by embedding the prediction of the neural network in the training objective.
  • To amortize this procedure over a potentially large number of evaluations, we also develop pre-trained HFL. To our knowledge, this is the first approach to regularization-based uncertainty that scales in the number of evaluations. An illustration of the set of methods considered in this paper is given in Figure 1.
The rest of the paper is organized as follows. In Section 2, we develop the Hessian-free Laplace method. Empirical evaluation is presented in Section 5 before discussing conclusions and future work in Section 6.
Figure 1: Summary of the regularization-based approach to uncertainty quantification that underlies Hessian-free Laplace. Step 1: fit or obtain an existing fit of a predictive model; step 2: fit another model using a differently regularized objective (from scratch or warm-started); step 3: the difference in predictions of these two models is the predictive variance. In Section 3, we present various types of regularization that make this strategy successful.
2 Bayesian Deep Learning with Hessian-Free Laplace
@@
Our starting point is the learning of parameters $\theta$ of a deep neural network $f_{\theta}$ as per Eq. 1 and 2. While the representational power and accuracy of the network may be high, in many applications it is crucial to also quantify the uncertainty of predictions, such as in autonomous vehicles, healthcare, risk-averse recommendations (Kendall & Gal, 2017; Leibig et al., 2017).
Bayesian probability theory provides a framework for analyzing uncertainty in deep learning, referred to as Bayesian deep learning (BDL). As for Bayesian inference in general, the key components are the prior distribution $p(\theta)$ and likelihood function $p(y\;|\;\theta,x)$ in a family of models indexed by $\theta$, and an i.i.d. dataset $\mathcal{D}=\left\{(x_{i},y_{i})\right\}_{i=1}^{n}$ assumed to be collected from a model in this family. The posterior distribution is then derived,
$p(\theta\;|\;\mathcal{D})=\frac{p(\mathcal{D}\;|\;\theta)p(\theta)}{\int p(\mathcal{D}\;|\;\theta)p(\theta)\mathrm{d}\theta},$ (4)
@@
and may be flexibly used as the density (or mass) for the expectation of any downstream prediction depending on $\theta$, e.g. expectation or variance of predicted mean, as part of a larger simulation involving fitted deep models, or interpreting the variance of weights for different layers of the network. The main evaluations of interest in BDL are the mean $\mathbb{E}_{p(\theta\;|\;\mathcal{D})}[f_{\theta}(x)]$ and (co)variance $\mathrm{Var}_{p(\theta\;|\;\mathcal{D})}[f_{\theta}(x)]$ of the prediction for a query vector (or set of query vectors) $x$. These will be the focus of our study in this paper.
Double Intractability
The steps in BDL discussed so far are doubly intractable in the following ways.
@@
  • Model Evidence For all but the most trivial networks, the integral in the denominator of Eq. 4, (the model evidence) entails a sum over high-dimensional $\theta$. Many techniques have been developed to approximate the posterior – the most common being MAP estimation, variational inference, and Markov chain Monte Carlo – and these vary in their scalability and applicability to deep neural nets (Blundell et al., 2015; Ritter et al., 2018). A key desideratum for approximate inference that is often overlooked is the extent to which it is compatible with existing deep learning frameworks and implementations, which may be supported by multiple years of model selection, tuning, and vast training budgets. The MAP point estimate is the most practical in this sense, and is the basis of the Laplace approximation.
  • Posterior Predictive The posterior predictive $p(y\;|\;\mathcal{D},x)$ is the expected predictive density, involving complex function $f_{\theta}$, under averages of the approximate posterior, meaning that the integral cannot be further simplified. Monte Carlo estimation using samples from the posterior is used to address this intractability.
@@
[…]
By inspection of Eq. 5, we see that,
$q(\theta)$ $\propto\exp\left\{-\frac{1}{2}(\theta-\hat{\theta})^{\top}P(\theta-\hat{\theta})\right\}$ (6)
$\implies q(\theta)$ $=\mathcal{N}(\hat{\theta},P^{-1}).$ (7)
If precision matrix $P$ can be calculated and inverted, then the first intractability of approximating the posterior is addressed. In addition, Eq. 7 also provides a closed-form solution to the model evidence, supporting model and hyperparameter selection (MacKay, 1992). We next discuss precision calculation in more detail.
@@
Eq. 7 depends on calculating the covariance matrix, i.e. inverting the precision matrix $P$ which can be decomposed as,
$P$ $=-\nabla\nabla_{\theta}\log p(\theta\;|\;\mathcal{D})$
$=\underbrace{-\nabla\nabla_{\theta}\log p(\theta)}_{\text{Gaussian prior}\implies\text{scalar matrix}}-\underbrace{\nabla\nabla_{\theta}\log p(\mathcal{D}\;|\;\theta)}_{\text{Hessian}\;H}.$ (8)
@@
The Hessian of the likelihood term may be further decomposed as,
$H$ $=\nabla\nabla_{\theta}\log p(y\;|\;f_{\theta}(x))$
$=\underbrace{\nabla_{f}\log p(y\;|\;f_{\theta}(x))}_{\text{residual}}{\nabla\nabla_{\theta}f_{\theta}(x)}+$
$\underbrace{\nabla\nabla_{f}\log p(y\;|\;f_{\theta}(x))}_{\text{observation precision}}{\nabla_{\theta}f_{\theta}(x)\nabla_{\theta}f_{\theta}(x)^{\top}}.$ (9)
Algorithm 1 Hessian-Free Laplace
  Warm start $\hat{\theta}^{(x,\lambda)}\leftarrow\hat{\theta}$
@@
     Follow stochastic gradient $\hat{\theta}^{(x,\lambda)}\leftarrow\hat{\theta}^{(x,\lambda)}+n\gamma(j)\nabla_{\theta}\tilde{\mathcal{L}}_{\theta}^{(x,\lambda)}(x_{i},y_{i})\;\;\;$ //Using Eq. 21.
@@
  until $\hat{\theta}^{(x,\lambda)}$ converges
In practice, the precision calculation of the Laplace approximation drops the first term – the network Hessian – in Eq. 9 due to computational constraints. This is known as generalized Gauss-Newton (GGN) and is proportional to the outer product of the network Jacobian (second term in Eq. 9). Beyond expediency, there are two arguments motivating GGN:
  1. 1.
    When fitting highly flexible models, as in BDL, the residual error is expected to be $\approx 0$.
  2. 2.
    The residual is a zero-mean random variable uncorrelated with the network Hessian, so the first term is approximately zero as the number of data points grows.
Immer et al. (2021) observed that when GGN is used in approximate inference, it assumes the network takes the form of a generalized linear model. In Laplace, this is equivalent to a local linearization of the network prediction around $\hat{\theta}$,
[…]
which experimentally results in more accurate predictions than the original network $f_{\hat{\theta}}$ when used in the Monte Carlo average for evaluation (Foong et al., 2019). However, the Hessian, even after simplifying with GGN, is still unwieldy or prohibitive to compute, store, and invert for large networks with high dimensional $\theta$.
3 Hessian-free Laplace (HFL)
[…]
${\mathcal{L}_{\theta}}^{(x,\lambda)}$ $=\sum_{i=1}^{n}\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\log p(\theta)+\lambda f_{\theta}(x),$ (12)
$\hat{\theta}^{(x,\lambda)}\in\arg_{\theta}\max{\mathcal{L}_{\theta}}^{(x,\lambda)}.$ (13)

### v2

@@ Abstract
Uncertainty quantification in deep learning is crucial for safe and reliable decision-making in downstream tasks. Existing methods quantify uncertainty at the last layer or other approximations of the network which may miss some sources of uncertainty in the model. To address this gap, we propose an uncertainty quantification method for large networks based on variation due to regularization. Essentially, predictions that are more (less) sensitive to the regularization of network parameters are less (more, respectively) certain. This principle can be implemented by deterministically tweaking the training loss during the fine-tuning phase and reflects confidence in the output as a function of all layers of the network. We show that regularization variation (RegVar) provides rigorous uncertainty estimates that, in the infinitesimal limit, exactly recover the Laplace approximation in Bayesian deep learning. We demonstrate its success in several deep learning architectures, showing it can scale tractably with the network size while maintaining or improving uncertainty quantification quality. Our experiments across multiple datasets show that RegVar not only identifies uncertain predictions effectively but also provides insights into the stability of learned representations.
@@ 1 INTRODUCTION
Figure 1: Summary of the regularization-based approach to uncertainty quantification that underlies RegVar. Step 1: fit a predictive model by MAP; step 2: fit another model using an infinitesimally regularized MAP; step 3: the difference between the predictions of these two models, normalized by the amount of regularization, is exactly the estimate of the predictive variance.
{mdframed}
// training step to update model parameters theta with training sample (x, y)// n: training data size, lam: scalar hyperparameterf = model(x)loss = cross_entropy(f, y) - log_prior(theta) / n + lam * f.abs().mean() / nloss.backward()optimizer.step()...//inference time for input x_testf_pred_var = (model(x_test) - f_pred_mean).abs() / lam
Figure 2: Example of fine-tuning a model using regularization variation (RegVar). The code required to extend existing fine-tuning to implement RegVar is given in blue. Any auto-differentiation package is sufficient, here we use PyTorch (Paszke et al., 2017).
@@
Bayesian neural networks, where a prior is placed on model weights, offer an opportunity to understand uncertainty and improve model robustness (Wilson, 2020; Papamarkou et al., 2024). However, exact computation of the model-weight posterior and predictive distributions is prohibitive given size and complexity. It is instead appealing to approximate it using the maximum a posteriori (MAP) solution and the curvature thereat. However, even this curvature, characterized by a Hessian, can be challenging to compute.
The Laplace approximation of the posterior distribution uses the (negative) Hessian of the log joint distribution at the MAP (Mackay, 1992). For high dimensional $\theta$ – common in deep neural networks – these steps are a prohibitive computational bottleneck even for an approximate posterior. One remedy is further approximation of the Hessian, but this may ignore important directions of curvature.
In this paper we propose using the principle of variation due to regularization (RegVar), which directly approximates the uncertainty of the predicted mean in the Laplace approximation without explicitly computing and inverting the Hessian (while still assuming it exists). Instead, one additional point estimate is required, the prediction-regularized MAP, derived from the network objective with a small amount of regularization added. Under the assumptions that the Laplace approximation is used, we find that the rescaled difference in network outputs given by the MAP and the prediction-regularized MAP recovers the Laplace variance.
Our contributions in this paper are as follows:
  • We develop a set of novel methods around regularization variation (RegVar) and formalize it as the predictive variance of linearized Laplace.
  • In RegVar, variance is derived from the change in the MAP prediction when adjusting the regularizer. The potential appeal of RegVar is the simplicity of the method (see Figure 1) and implementation (see Figure 2 for an example).
  • We show that RegVar scales up to large networks and compare it to several other popular Bayesian deep learning approaches.
The rest of the paper is organized as follows. In Section 2, we discuss related work. In Section 3, we set up the key aspects of the Laplace approximation and connect it to regularization variation in Section 4. Empirical evaluation is explored in Section 5 before discussing conclusions and future work in Section 6.
2 RELATED WORK
The Laplace approximation was first formulated in BDL for small-scale neural networks (MacKay, 1992). Since then, larger architectures have necessitated Hessian approximations, including Gauss-Newton (Foresee and Hagan, 1997) and Kronecker factorization that places a block diagonal structure on the covariate in line with assumed independence between the layers of a network (Martens and Grosse, 2015).
In recent years, with the aforementioned scalable approximations and renewed appreciation of the advantages of Laplace as a post-hoc uncertainty method (Daxberger et al., 2021), a number of new directions were addressed, including relating neural network loss functions to Gaussian process inference (Khan et al., 2019) and exploring the advantages of locally linearized Laplace in out-of-distribution evaluation (Foong et al., 2019). Underfitting in Laplace (Lawrence, 2001) is mitigated in the GGN approximation by switching to the generalized linear model implied by the local linearization (Immer et al., 2021). Further scaling with linearized Laplace applies neural tangent kernels (Deng et al., 2022) and variational approximation (Ortega et al., 2023). For large models, Laplace can be applied to a low-rank adaptor Yang et al. (2024).
Beyond Laplace-based methods, a number of approaches have been proposed for scalable Bayesian inference in neural networks. Given the breadth of this topic, we give a non-exhaustive account (see Papamarkou et al., 2024 for further discussion). Variational inference approximates the posterior with a parameterized distribution, enabling efficient optimization via stochastic gradient descent (Graves, 2011; Blundell et al., 2015; Shen et al., 2024). Stochastic gradient Langevin dynamics and related approaches that take samples from the optimization trajectory of a neural network use gradient noise to approximate Bayesian updates without explicit posterior evaluation (Welling and Teh, 2011; Mandt et al., 2016; Maddox et al., 2019). Dropout-based uncertainty estimation interprets dropout training as an approximate Bayesian inference method (Gal and Ghahramani, 2016). Ensembles of independently trained networks use multiple predictions to quantify the uncertainty of black-box models (Lakshminarayanan et al., 2017; Osband and Van Roy, 2015).
Owing to the ubiquity of the Hessian of the objective and its computational demands, approaches to avoiding its direct evaluation are developed in areas of research distinct of Laplace. In particular, Hessian-free optimization uses the conjugate gradient method to iteratively update a matrix-vector product as part of an inner loop in Newton’s method second-order optimization (Martens and Sutskever, 2011; Pearlmutter, 1994). Finally, Kallus and McInerney (2022) develop a frequentist approach establishing that infinitesimal regularization of the total log likelihood yields a variance estimate that is asymptotically equivalent to the delta method. While avoiding full Fisher matrix computation, this approach requires evaluating a perturbed estimator per test point and output dimension, leading to computational costs that scale accordingly. It is also a frequentist perspective, in contrast to the Bayesian approach developed in this paper.
3 BACKGROUND
@@
$\hat{\theta}$ $\in\arg_{\theta}\max\mathcal{L}_{\theta}$ (1)
@@
for some observation likelihood $p(y\;|\;\cdot)$ and prior $p(\theta)$. This is equivalent to optimizing a loss (e.g., squared loss for Gaussian likelihood or cross-entropy for categorical likelihood) regularized by $\log p(\theta)$ and means that computing the MAP of a Bayesian neural network is akin to fitting a vanilla neural network.
@@
Our starting point is the learning of parameters $\theta$ of a deep neural network $f_{\theta}$ as per Eq. 1 and 2. While the representational power and accuracy of the network may be high, in many applications it is crucial to also quantify the uncertainty of predictions, such as in autonomous vehicles, healthcare, or recommendations (Kendall and Gal, 2017; Leibig et al., 2017).
@@
Bayesian deep learning (BDL) provides a framework for analyzing uncertainty in deep learning (Papamarkou et al., 2024). The key components are the prior distribution $p(\theta)$ and likelihood function $p(y\;|\;\theta,x)$ in a family of models indexed by $\theta$, and an i.i.d. dataset $\mathcal{D}=\left\{(x_{i},y_{i})\right\}_{i=1}^{n}$ assumed to be collected from a model in this family. The posterior distribution $p(\theta\;|\;\mathcal{D})$ may be flexibly used as the density (or mass) for the expectation of any downstream prediction depending on $\theta$, e.g. expectation or variance of predicted mean, as part of a larger simulation involving fitted deep models, or interpreting the variance of weights for different layers of the network. The main evaluations of interest in BDL are the mean $\mathbb{E}_{p(\theta\;|\;\mathcal{D})}[f_{\theta}(x)]$ and (co)variance $\mathrm{Var}_{p(\theta\;|\;\mathcal{D})}[f_{\theta}(x)]$ of the prediction for a query vector (or set of query vectors) $x$. These will be the focus of our study in this paper.
@@
BDL is doubly intractable. First, the normalizing factor in the Bayesian posterior is intractable for all but the most trivial networks as it entails a sum over high-dimensional $\theta$. Many techniques have been developed to approximate the posterior – the most common being MAP estimation, variational inference, and Markov chain Monte Carlo – and these vary in their scalability and applicability to deep neural nets (Blundell et al., 2015; Ritter et al., 2018). A key desideratum for approximate inference that is often overlooked is the extent to which it is compatible with existing deep learning frameworks and implementations, which may be supported by years of model selection, tuning, and vast training budgets. The MAP point estimate is the most practical in this sense, and is the basis of the Laplace approximation. Second, even with an approximate posterior to hand, the posterior predictive $p(y\;|\;\mathcal{D},x)$ depends on $f_{\theta}$, meaning that the integral cannot be further simplified. Monte Carlo estimation using samples from the posterior is used to address this intractability.
@@
where $P:=-\nabla\nabla_{\theta}\log p(\theta\;|\;\mathcal{D})$. The first derivative does not appear in Eq. 3 because $\hat{\theta}$ is defined as a local optimum of the posterior, where it exists.
Noticing that Eq. 3 takes the form of a quadratic in $\theta$, we see that $q(\theta)=\mathcal{N}(\hat{\theta},P^{-1})$. If precision matrix $P$ can be calculated and inverted, then the first intractability of approximating the posterior is addressed. In addition, the normal form of $q$ also provides a closed-form solution to the model evidence, supporting model and hyperparameter selection (MacKay, 1992). We next discuss precision calculation in more detail.
@@
Distribution $q(\theta)$ requires the covariance matrix which entails calculating and inverting the precision matrix $P$ decomposed as,
$P$ $=-\nabla\nabla_{\theta}\log p(\theta\mid\mathcal{D})$
$=\underbrace{-\nabla\nabla_{\theta}\log p(\theta)}_{\text{Gaussian prior}\implies\text{scalar matrix}}\quad-\underbrace{\nabla\nabla_{\theta}\log p(\mathcal{D}\mid\theta)}_{\text{Hessian}\;H}.$ (4)
@@
The Hessian of the likelihood term in Eq. 4 may be further decomposed as,
$H$ $=\underbrace{\nabla_{f}\log p(y\mid f_{\theta}(x))}_{\text{residual}}{\nabla\nabla_{\theta}f_{\theta}(x)}+$
$\underbrace{\nabla\nabla_{f}\log p(y\mid f_{\theta}(x))}_{\text{observation precision}}\quad\nabla_{\theta}f_{\theta}(x)\nabla_{\theta}f_{\theta}(x)^{\top}.$ (5)
In practice, the precision calculation of the Laplace approximation drops the first term – the network Hessian – in Eq. 5 due to computational constraints. This is known as generalized Gauss-Newton (GGN) and is proportional to the outer product of the network Jacobian (second term in Eq. 5). When GGN is used in approximate inference, it assumes the network takes the form of a generalized linear model. In Laplace, this is equivalent to a local linearization of the network prediction around $\hat{\theta}$,
$\tilde{f}_{\theta}(x)=f_{\hat{\theta}}(x)+\nabla_{\theta}f_{\theta}(x)\biggr{\rvert}_{\hat{\theta}}^{\top}(\theta-\hat{\theta}),$ (6)
which experimentally results in more accurate predictions than the original network $f_{\hat{\theta}}$ when used in the Monte Carlo average for evaluation (Foong et al., 2019; Immer et al., 2021). However, the Hessian, even after simplifying with GGN, is still unwieldy or prohibitive to compute, store, and invert for large networks with high dimensional $\theta$.
4 REGULARIZATION VARIATION (RegVar)
where $\nabla_{\theta}f_{\hat{\theta}}:=\nabla_{\theta}f_{\theta}\biggr{\rvert}_{\hat{\theta}}$ is used for more compact notation. This is an instance of the Bayesian delta method (Wasserman, 2006), differing from the classic delta method by the inclusion of the prior term in $P$ in Eq. 4.
${\mathcal{L}_{\theta}}^{(f(x),\lambda)}$ $=\sum_{i=1}^{n}\log p(y_{i}\;|\;f_{\theta}(x_{i}))+\log p(\theta)+\lambda f_{\theta}(x),$ (8)
$\hat{\theta}^{(f(x),\lambda)}\in\arg_{\theta}\max{\mathcal{L}_{\theta}}^{(f(x),\lambda)}.$ (9)
Theorem 4.1.
$\nabla_{\lambda}f_{\hat{\theta}^{(f(x),\lambda)}}(x)\biggr{\rvert}_{\lambda=0}=\nabla_{\theta}f_{\hat{\theta}}(x)^{\top}P^{-1}\nabla_{\theta}f_{\hat{\theta}}(x),$ (10)
assuming $\hat{\theta}^{(f(x),\lambda)}$ and $P^{-1}$ exist, and that ${\mathcal{L}_{\theta}}^{(f(x),\lambda)}$ is continuously differentiable w.r.t. $\theta$.
Proof.
See Appendix A. ∎
A parallel frequentist approach is known as the implicit delta method (Kallus and McInerney, 2022). The relationship in Eq. 10 can also be understood as an application of the implicit function theorem (IFT) by Cauchy, e.g. see (Lorraine et al., 2020) for a modern treatment of IFT.
Algorithm 1 Regularization Variation (RegVar)
  Input: network $f_{\theta}$, fitted MAP $\hat{\theta}$, training data $\mathcal{D}$ of size $n$, evaluation inputs $\{\hat{x}_{i=1}^{m}\}$, scalar $\lambda$, step size schedule $\gamma(\cdot)$
  Warm start $\hat{\theta}^{(\hat{r},\lambda)}\leftarrow\hat{\theta}$
@@ 4 REGULARIZATION VARIATION (RegVar)
     Follow gradient $\hat{\theta}^{(\hat{r},\lambda)}\leftarrow\hat{\theta}^{(\hat{r},\lambda)}+n\gamma(j)\nabla_{\theta}\mathcal{L}_{i}\;\;\;$ // see Eq. 15
@@
  until $\hat{\theta}^{(\hat{r},\lambda)}$ converges
  Return: predictive variance function $\hat{\sigma}_{f}^{2}(x):=\frac{1}{\lambda}|f_{\hat{\theta}^{(\hat{r},\lambda)}}(x)-f_{\hat{\theta}}(x)|$
$\frac{1}{\lambda}(f_{\hat{\theta}^{(f(x),\lambda)}}(x)-f_{\hat{\theta}}(x))\underset{\lambda\to 0}{\rightarrow}\nabla_{\lambda}f_{\hat{\theta}^{(f(x),\lambda)}}(x)\biggr{\rvert}_{\lambda=0}.$ (11)
It is natural to consider auto-differentiation to calculate the LHS of Eq. 10. From the fact that the optimum $\hat{\theta}^{(f(x),\lambda)}$ also changes with $\lambda$ it can be seen that auto-differentiation must deal with complex operations on optima and is unlikely to lead to a computational advantage unless used in tandem with additional work-arounds.
Putting it together, the regularization variation (RegVar) approximation is,
$\tilde{f}_{\theta}\sim\mathcal{N}\left(f_{\hat{\theta}}(x),\frac{1}{\lambda}(f_{\hat{\theta}^{(f(x),\lambda)}}(x)-f_{\hat{\theta}}(x))\right),$ (12)
and differs from linearized Laplace (Eq. 7) by the lack of an explicit Hessian. It is apparent by this comparison that RegVar trades the computational and storage expense of dealing with the Hessian for that of a point-wise approximation for each query $x$. This operation may be amortized into the network, which we discuss next.
4.1 Amortized Regularizer
RegVar targets the same predictive variance-covariance as the full linearized Laplace approximation. However, when evaluating the uncertainty of many predictions, the requirement to optimize to the $f$-objective specific to each test point is a bottleneck. In order to scale RegVar in the number of test points, we develop an amortized regularization approach that requires only a single model to represent the $f$-regularized parameters for a set of evaluation points.
Amortization is formulated as follows. Find the regularization term $\hat{r}$ for a finite set of evaluation inputs ${\bf\hat{x}}=\{\hat{x}_{1:m}\}$ that minimizes the average squared L2 norm error of the corresponding $f$-regularized parameters,
$\hat{r}=\arg_{r}\min\sum_{i=1}^{m}\|\hat{\theta}^{(f(x_{i}),\lambda)}-\hat{\theta}^{(r,\lambda)}\|_{2}^{2}.$ (13)
Eq. 13 facilitates the calculation of Eq. 12 using $\hat{\theta}^{(\hat{r},\lambda)}$ instead of $\hat{\theta}^{(f(x),\lambda)}$ for arbitrary $x$ in the domain of $f$.